# The Ranking-Selberg integral on ${\bf GSp}(2)$ for square free levels

### v1

@@ 4.4. Bessel models
It is known that $\dim_{{\mathbb{C}}}(V_{\pi}^{*})^{T_{\theta},\Lambda}\leqslant 1$ ([27], [28]). We say that $\pi_{p}$ has a local $(T_{\theta},\Lambda_{p})$-Bessel model if $(V_{\pi}^{*})^{T_{\theta},\Lambda}\not=(0)$; when this is the case, the space of functions of the form $g\mapsto\ell(\pi_{p}(g)\xi)$ with $\xi\in V_{\pi_{p}}$ is independent of $\ell\in(V_{\pi_{p}}^{*})^{T_{\theta},\Lambda_{p}}-(0)$; this space is denoted by ${\mathscr{B}}(T_{\theta},\Lambda_{p})[\pi_{p}]$ and is called the local $(T_{\theta},\Lambda_{p})$-Bessel model of $\pi_{p}$. When $\pi_{p}$ is spherical, it is known that $\pi_{p}$ has a local $(T_{\theta},\Lambda_{p})$-Bessel model, and the space ${\mathscr{B}}(T_{\theta},\Lambda_{p})[\pi_{p}]$ contains a unique ${\bf K}_{p}$-invariant function $B_{\pi_{p}}$ such that $B_{\pi_{p}}^{0}(1_{4})=1$ ([43, Theorem 2-I], [6]). Let $\xi_{p}^{0}$ be the non-zero ${\bf K}_{p}$-fixed vector in $V_{\pi_{p}}$; then there exists a unique element $\ell_{\pi_{p}}^{0}\in(V_{\pi_{p}}^{*})^{T_{\theta},\Lambda_{p}}-(0)$ such that
$\ell_{\pi_{p}}^{0}(\pi_{p}(g)\xi_{\pi_{p}}^{0})=B_{\pi_{p}}^{0}(g),\quad g\in\mathsf{G}({\mathbb{Q}}_{p}).$
The pair $(\ell_{\pi_{p}}^{0},\xi_{\pi_{p}}^{0})$ is referred to as the unramified $(T_{\theta},\Lambda_{p})$-Bessel datum for $\pi_{p}$.
Let $\pi\cong\otimes_{p\leqslant\infty}\pi_{p}\in\Pi_{\rm cusp}^{(T_{\theta},\Lambda)}(\lambda,N)$ (see §4.3). Let $N_{\pi}$ be the product of $p\mid N$ such that $\pi_{p}$ is not spherical.
@@
A system $\{(\ell_{p},\xi_{p})\}_{p<\infty}$ with $\ell_{p}\in(V_{\pi_{p}}^{*})^{T_{\theta},\Lambda_{p}}-(0)$, $\xi_{p}\in V_{\pi_{p}}^{{\bf K}_{0}(N_{\pi}{\mathbb{Z}}_{p})}$ is called a $(T_{\theta},\Lambda)$-Bessel data for $\pi$ if $(\ell_{p},\xi_{p})=(\ell_{\pi_{p}}^{0},\xi_{\pi_{p}}^{0})$ for $(p,N_{\pi})=1$ and $\ell_{p}(\xi_{p})=1$ for all $p<\infty$.
By [32, Theorem 2 8.2 and 9.3], a $(T_{\theta},\Lambda)$-Bessel data exists for our $\pi$. Once a $(T_{\theta},\Lambda)$-Bessel data $\{(\ell_{p},\xi_{p})\}_{p<\infty}$ is fixed, one can define $\varphi_{\pi,\varrho}^{0}\in V_{\pi}^{{\bf K}_{0}(N_{\pi})}[\varrho]$ to be the $V_{\varrho}$-valued cusp form such that for any $v\in V_{\varrho}$, the function $(\varphi_{\pi,\varrho}^{0}(g)\mid v)_{\varrho}$ in $V_{\pi}\cong\otimes_{p\leqslant\infty}V_{\pi_{p}}$ corresponds to the pure tensor $\bar{v}\otimes(\otimes_{p<\infty}\xi_{p})$, where $\bar{v}=1\otimes v\in\overline{V_{\varrho}}\hookrightarrow V_{\pi_{\infty}}$. A particular choice of $\{(\ell_{p},\xi_{p})\}$ will be made in §6.8 so that $\varphi_{\pi,\varrho}^{0}$ corresponds to a newform on the arithmetic quotient $\Gamma_{0}(N)\backslash{\mathfrak{h}}_{2}$ in the sense of [7, §3.2].
@@
(4.20) $\langle E(\phi,s,\Lambda,\mu),R(b_{\rm{\bf{f}}}b_{\mathbb{R}}^{\theta})\overline{\varphi_{v}}\rangle=\tfrac{\sqrt{|D|}}{2}Z_{v}^{(\infty)}(\phi_{\infty},\varphi^{0}_{\pi};s,\Lambda)\times\prod_{p\leqslant\infty}Z_{p}(\phi_{p},\overline{B_{v_{p}}};s,\mu_{p};b_{p}),$
@@
$Z_{v}^{(\infty)}(\phi_{\infty},\varphi^{0}_{\pi};s,\Lambda)$ $:=\int_{{\bf K}_{\infty}^{\#}}\int_{{\mathbb{R}}^{\times}}|a|_{\mathbb{R}}^{s-1}\,f_{\phi_{\infty}}^{(s,1,1)}(k_{\infty}^{\#})(v\mid B^{T_{\theta},\Lambda}(\varphi^{0}_{\pi,\varrho};\mathsf{m}(a1_{2},a)\iota_{\theta}(k^{\#})b_{\mathbb{R}}^{\theta}))_{\varrho}\,{\rm{d}}^{\times}a\,{\rm{d}}k^{\#}_{\infty},$
$Z_{p}(\phi_{p},\overline{B_{v_{p}}};s,\mu_{p};b_{p})$ $:=\int_{{\mathbb{Q}}_{p}^{\times}}\int_{{\bf K}_{p}^{\#}}f_{\phi_{p}}^{(s,\Lambda_{p},\mu_{p})}(k_{p}^{\#})\mu_{p}(a)|a|_{p}^{s-1}\,\overline{B_{v_{p}}}(\mathsf{m}(a_{p}1_{2},a_{p})\iota_{\theta}(k_{p}^{\#})b_{p})\,{\rm{d}}^{\times}a_{p}\,{\rm{d}}k^{\#}_{p}$
with
$B_{v_{p}}(g_{p}):=\ell_{p}(\pi_{p}(g_{p})v_{p}),\quad g_{p}\in\mathsf{G}({\mathbb{Q}}_{p}).$
@@ 4.5. Gamma factor and global Bessel period
Recall the point $b_{\mathbb{R}}^{\theta}$ in (2.14) and the vector $v_{\varrho}^{\theta}\in V_{\varrho}$ in (4.16). Since $\nu(b_{{\mathbb{R}}}^{\theta})>0$, the point $\mathsf{m}(a1_{2},a)b_{{\mathbb{R}}}^{\theta}$ with $a\in{\mathbb{R}}^{\times}$ belongs to $\mathsf{G}({\mathbb{R}})^{0}$ if and only if $a>0$. Thus, by (3.6) and (4.6), we have $B^{T_{\theta},\Lambda}(\varphi_{\pi}^{0};\mathsf{m}(a1_{2},a)b_{\mathbb{R}}^{\theta})=0$ for $a<0$. By Lemma 4.4 and (4.16),
@@
(4.22) $w_{D}^{-1}\overline{R(\varphi^{0}_{\pi},E,\Lambda)}\int_{{\bf K}^{\#}_{\infty}}\int_{0}^{\infty}|a|_{\mathbb{R}}^{s-1}\,f_{\phi_{\infty}}^{(s,1,1)}(k_{\infty}^{\#})(v\mid B^{T_{\theta}}_{\varrho}(\mathsf{m}(a1_{2},a)\iota_{\theta}(k_{\infty}^{\#})b_{\mathbb{R}}^{\theta}))_{\varrho}\,{\rm{d}}^{\times}a\,{\rm{d}}k^{\#}_{\infty}.$
@@
$\mathsf{f}_{\varrho}(u):=(d+1)\frac{(\varrho(C\bar{u}C^{-1})\mathsf{v}_{\varrho}^{0}\mid\mathsf{v}_{\varrho}^{0})_{\varrho}}{(\mathsf{v}_{\varrho}^{0}\mid\mathsf{v}_{\varrho}^{0})_{\varrho}},\qquad u\in{\bf K}^{\#}_{\infty}.$
with $C:=\frac{1}{\sqrt{2}}\left[\begin{smallmatrix}1&1\\ i&-i\end{smallmatrix}\right]$. If $u=\left[\begin{smallmatrix}a&-\bar{b}\\ b&\bar{a}\end{smallmatrix}\right]\in{\bf SU}(2)$ with $a=a^{\prime}+ia^{\prime\prime},\,b=b^{\prime}+ib^{\prime\prime}$ and $A,B\in{\bf Mat}_{2}({\mathbb{R}})$ are defined as in (2.16), so that $(b_{\mathbb{R}}^{\theta})^{-1}\iota_{\theta}(u)(b_{\mathbb{R}}^{\theta}=k_{\infty}(A+iB)$ (Lemma 2.2), then a computation reveals $C\bar{u}C^{-1}=A-iB$. Thus, the automorphism $u\mapsto C\bar{u}C^{-1}$ of ${\bf SU}(2)$ brings the subgroup ${\mathsf{B}}^{\#}({\mathbb{R}})\cap({\bf K}^{\#}_{\infty})^{0}=\{\left[\begin{smallmatrix}a&0\\ 0&a^{-1}\end{smallmatrix}\right]\mid a\in{\mathbb{C}}^{1}\}$ to the subgroup ${\bf SO}(2)$, by which $\mathsf{v}_{\varrho}^{0}$ is fixed. Hence, $\mathsf{f}_{\varrho}$ is left ${\mathsf{B}}^{\#}({\mathbb{R}})\cap({\bf K}^{\#}_{\infty})^{0}$-invariant. Since $\mathsf{v}_{\varrho}^{0}$ is also fixed by the element $C\delta C^{-1}=\delta$, the function $\mathsf{f}_{\varrho}$ is left ${\mathsf{B}}^{\#}({\mathbb{R}})\cap{\bf K}^{\#}_{\infty}$-invariant. Thus, there exists a unique element $\mathsf{f}^{(s)}_{\varrho}$ of ${\mathcal{V}}^{\#}(s,1,1)$ such that $\mathsf{f}_{\varrho}^{(s)}|{\bf K}_{\infty}^{\#}=\mathsf{f}_{\varrho}$. Define $\phi_{\infty}\in{\mathcal{S}}({\mathbb{C}}^{2})$ by
@@
(4.24) $f_{\phi_{\infty}}^{(s,1,1)}=|D|^{\frac{s+1}{2}}(-1)^{\frac{l_{1}-l_{2}}{2}}{\Gamma_{{\mathbb{C}}}\left(s+\tfrac{l_{1}-l_{2}}{2}+1\right)}\,\mathsf{f}_{\varrho}^{(s)}.$
@@
(4.25) $Z^{(\infty)}_{v}(\phi_{\infty},\varphi^{0}_{\pi};s,\Lambda)$ $=w_{D}^{-1}\overline{R(\varphi^{0}_{\pi},E,\Lambda)}(v\mid v_{\varrho}^{0})_{\varrho}(-1)^{l_{1}}|D|^{\frac{1}{2}(1-\frac{l_{1}+l_{2}}{2})}{L\left(s+\tfrac{1}{2},\pi_{\infty}\right)}.$
The formula in [47, 7.23 Lemma] yields
@@
(4.27) $f_{\widehat{\phi_{\infty}}}^{(-s,1,1)}=\tfrac{{|D|}}{4}(-1)^{\frac{l_{1}-l_{2}}{2}}f_{\phi_{\infty}}^{(-s,1,1)}.$
@@ 4.6. The spinor $L$-function and its functional equation
Let $\pi\cong\otimes_{v}\pi_{v}$ be a cuspidal automorphic representation of $\mathsf{G}({\mathbb{A}})$ with the trivial central character: then, $\bar{\pi}_{p}\cong\pi_{p}^{\vee}\cong\pi_{p}$ for all $p<\infty$ by [44, Proposition 2.3]. The twist $\pi_{\mu}$ of $\pi$ by an idele class character $\mu:{\mathbb{A}}^{\times}/{\mathbb{Q}}^{\times}\longrightarrow{\mathbb{C}}^{1}$ is defined on the space $V_{\pi}$ of $\pi$ as $\pi_{\mu}(g):=\pi(g)\cdot(\mu\circ\nu)(g)$ for $g\in\mathsf{G}({\mathbb{A}})$, so that the central character of $\pi_{\mu}$ is $\mu^{2}$. Let $\pi\in\Pi_{\rm cusp}^{(T_{\theta},\Lambda)}(\lambda,N)$ and $\mu$ be as in §4.3, so that $\lambda=(l_{1},l_{2})\in{\mathbb{Z}}^{2}$, $l_{1}\equiv l_{2}\pmod{2}$, $l_{1}\geqslant l_{2}>2$ and the ramification loci of $\pi$ and $\mu$ are disjoint. We define the spinor $L$-function of $\pi$ twisted by $\mu$ as the Euler product
@@
with $L(s,(\pi_{\mu})_{p})$ the local $L$-factor listed in table (6.15). The completed $L$-function is defined as
@@
as expected. Moreover $\widehat{L}(s,\pi,\mu)$ is holomorphic except for possible simple poles at $s=3/2,-1/2$, which does not occur when $\mu$ is non-trivial. By (4.26), $M(s)\mathsf{f}_{\varrho}^{(s)}$ has a simple zero at $s=1$ if $l_{1}>l_{2}$, which in turn implies the holomorphy at $s=1$ of the global intertwinign operator applied to the section $f_{\phi}^{(s)}$ as well as the Eisenstein series $E(\phi,s,\Lambda,\mu)$ for $\phi_{\infty}$ as above. Hence, by (4.20) with an appropriate $\phi$, $\widehat{L}(s,\pi,\mu)$ is holomorphic at $s=3/2$ when $l_{1}>l_{2}$. If $\mu$ is real valued, then $G(\mu)=\sqrt{M}$ and $\varepsilon(1/2,\pi,\mu)=(-1)^{l_{2}}$, so that $L(1/2,\pi,\mu)=0$ unless $l_{2}$ is even.

### v2

@@ 4.4. Bessel models
It is known that $\dim_{{\mathbb{C}}}(V_{\pi}^{*})^{T_{\theta},\Lambda}\leqslant 1$ ([28], [29]). We say that $\pi_{p}$ has a local $(T_{\theta},\Lambda_{p})$-Bessel model if $(V_{\pi}^{*})^{T_{\theta},\Lambda}\not=(0)$; when this is the case, the space of functions of the form $g\mapsto\ell(\pi_{p}(g)\xi)$ with $\xi\in V_{\pi_{p}}$ is independent of $\ell\in(V_{\pi_{p}}^{*})^{T_{\theta},\Lambda_{p}}-(0)$; this space is denoted by ${\mathscr{B}}(T_{\theta},\Lambda_{p})[\pi_{p}]$ and is called the local $(T_{\theta},\Lambda_{p})$-Bessel model of $\pi_{p}$. When $\pi_{p}$ is spherical, it is known that $\pi_{p}$ has a local $(T_{\theta},\Lambda_{p})$-Bessel model, and the space ${\mathscr{B}}(T_{\theta},\Lambda_{p})[\pi_{p}]$ contains a unique ${\bf K}_{p}$-invariant function $B_{\pi_{p}}$ such that $B_{\pi_{p}}^{0}(1_{4})=1$ ([44, Theorem 2-I], [6]). Let $\xi_{p}^{0}$ be the non-zero ${\bf K}_{p}$-fixed vector in $V_{\pi_{p}}$; then there exists a unique element $\ell_{\pi_{p}}^{0}\in(V_{\pi_{p}}^{*})^{T_{\theta},\Lambda_{p}}-(0)$ such that $\ell_{\pi_{p}}^{0}(\pi_{p}(g)\xi_{\pi_{p}}^{0})=B_{\pi_{p}}^{0}(g)$ for all $g\in\mathsf{G}({\mathbb{Q}}_{p})$. The pair $(\ell_{\pi_{p}}^{0},\xi_{\pi_{p}}^{0})$ is referred to as the unramified $(T_{\theta},\Lambda_{p})$-Bessel datum for $\pi_{p}$.
Let $\pi\cong\otimes_{p\leqslant\infty}\pi_{p}\in\Pi_{\rm cusp}^{(T_{\theta},\Lambda)}(\lambda,N)$ (see §4.3).
@@
A system $\{(\ell_{p},\xi_{p})\}_{p<\infty}$ with $\ell_{p}\in(V_{\pi_{p}}^{*})^{T_{\theta},\Lambda_{p}}-(0)$, $\xi_{p}\in V_{\pi_{p}}^{{\bf K}_{0}(N_{\pi}{\mathbb{Z}}_{p})}$ is called a $(T_{\theta},\Lambda)$-Bessel data for $\pi$ if $(\ell_{p},\xi_{p})=(\ell_{\pi_{p}}^{0},\xi_{\pi_{p}}^{0})$ for $p\nmid N_{\pi}$ and $\ell_{p}(\xi_{p})=1$ for all $p<\infty$.
By [33, Theorem 2 8.2 and 9.3], a $(T_{\theta},\Lambda)$-Bessel data exists for our $\pi$. Once a $(T_{\theta},\Lambda)$-Bessel data $\{(\ell_{p},\xi_{p})\}_{p<\infty}$ is fixed, one can define $\varphi_{\pi,\varrho}^{0}\in V_{\pi}^{{\bf K}_{0}(N_{\pi})}[\varrho]$ to be the $V_{\varrho}$-valued cusp form such that for any $v\in V_{\varrho}$, the function $(\varphi_{\pi,\varrho}^{0}(g)\mid v)_{\varrho}$ in $V_{\pi}\cong\otimes_{p\leqslant\infty}V_{\pi_{p}}$ corresponds to the pure tensor $\bar{v}\otimes(\otimes_{p<\infty}\xi_{p})$, where $\bar{v}=1\otimes v\in\overline{V_{\varrho}}\hookrightarrow V_{\pi_{\infty}}$. A particular choice of $\{(\ell_{p},\xi_{p})\}$ will be made in §6.9 so that $\varphi_{\pi,\varrho}^{0}$ corresponds to a newform on the arithmetic quotient $\Gamma_{0}(N)\backslash{\mathfrak{h}}_{2}$ in the sense of [7, §3.2].
@@
(4.20) $\langle E(\phi,s,\Lambda,\mu),R(b_{\rm{\bf{f}}}b_{\mathbb{R}}^{\theta})\overline{\varphi_{v}}\rangle=\tfrac{\sqrt{|D|}}{2}Z_{v}^{(\infty)}(\phi_{\infty},\varphi^{0}_{\pi};s,\mu_{\infty},\Lambda)\times\prod_{p\leqslant\infty}Z_{p}(\phi_{p},\overline{B_{v_{p}}};s,\mu_{p};b_{p}),$
@@
$Z_{v}^{(\infty)}(\phi_{\infty},\varphi^{0}_{\pi};s,\mu_{\infty},\Lambda)$ $:=\int_{{\bf K}_{\infty}^{\#}}\int_{{\mathbb{R}}^{\times}}f_{\phi_{\infty}}^{(s,1,\mu_{\infty})}(k_{\infty}^{\#})\mu_{\infty}(a)|a|_{\mathbb{R}}^{s-1}$
×$\qquad\times(v\mid B^{T_{\theta},\Lambda}(\varphi^{0}_{\pi,\varrho};\mathsf{m}(a1_{2},a)\iota_{\theta}(k^{\#})b_{\mathbb{R}}^{\theta}))_{\varrho}\,{\rm{d}}^{\times}a\,{\rm{d}}k^{\#}_{\infty},$
$Z_{p}(\phi_{p},\overline{B_{v_{p}}};s,\mu_{p};b_{p})$ $:=\int_{{\bf K}_{p}^{\#}}\int_{{\mathbb{Q}}_{p}^{\times}}f_{\phi_{p}}^{(s,\Lambda_{p},\mu_{p})}(k_{p}^{\#})\mu_{p}(a)|a|_{p}^{s-1}\,\overline{B_{v_{p}}}(\mathsf{m}(a_{p}1_{2},a_{p})\iota_{\theta}(k_{p}^{\#})b_{p})\,{\rm{d}}^{\times}a_{p}\,{\rm{d}}k^{\#}_{p}$
with $B_{v_{p}}(g_{p}):=\ell_{p}(\pi_{p}(g_{p})v_{p})$ for $g_{p}\in\mathsf{G}({\mathbb{Q}}_{p})$.
@@ 4.5. Gamma factor and global Bessel period
Recall the point $b_{\mathbb{R}}^{\theta}$ in (2.14) and the vector $v_{\varrho}^{\theta}\in V_{\varrho}$ in (4.16). Since $\nu(b_{{\mathbb{R}}}^{\theta})>0$, the point $\mathsf{m}(a1_{2},a)\iota_{\theta}(k^{\#}_{\infty})b_{{\mathbb{R}}}^{\theta}$ with $a\in{\mathbb{R}}^{\times}$ and $k_{\infty}^{\#}\in{\bf K}_{\infty}^{\#}$ belongs to $\mathsf{G}({\mathbb{R}})^{0}$ if and only if ${\rm sgn}(a){\rm sgn}(k_{\infty}^{\#})=+1$, where ${\rm sgn}(k_{\infty}^{\#}):={\rm sgn}(\nu(\iota_{\theta}(k_{\infty}^{\#}))$. Thus, by (3.7) and (4.6), we have $B^{T_{\theta},\Lambda}(\varphi_{\pi}^{0};\mathsf{m}(a1_{2},a)\iota_{\theta}(k_{\infty}^{\#})b_{\mathbb{R}}^{\theta})=0$ unless ${\rm sgn}(a){\rm sgn}(k_{\infty}^{\#})=+1$. By Lemma 4.4 and (4.16),
@@
(4.22) $2w_{D}^{-1}\overline{R(\varphi^{0}_{\pi},E,\Lambda)}\int_{({\bf K}^{\#}_{\infty})^{0}}\int_{0}^{\infty}|a|_{\mathbb{R}}^{s-1}\,f_{\phi_{\infty}}^{(s,1,\mu_{\infty})}(k_{\infty}^{\#})(v\mid B^{T_{\theta}}_{\varrho}(\mathsf{m}(a1_{2},a)\iota_{\theta}(k_{\infty}^{\#})b_{\mathbb{R}}^{\theta}))_{\varrho}\,{\rm{d}}^{\times}a\,{\rm{d}}k^{\#}_{\infty}.$
@@
$\mathsf{f}_{\varrho}(u):=(d+1)\frac{(\varrho(C\bar{u}C^{-1})\mathsf{v}_{\varrho}^{0}\mid\mathsf{v}_{\varrho}^{0})_{\varrho}}{(\mathsf{v}_{\varrho}^{0}\mid\mathsf{v}_{\varrho}^{0})_{\varrho}},\qquad u\in({\bf K}^{\#}_{\infty})^{0}$
with $C:=\frac{1}{\sqrt{2}}\left[\begin{smallmatrix}1&1\\ i&-i\end{smallmatrix}\right]$. If $u=\left[\begin{smallmatrix}a&-\bar{b}\\ b&\bar{a}\end{smallmatrix}\right]\in({\bf K}_{\infty}^{\#})^{0}={\bf SU}(2)$ with $a=a^{\prime}+ia^{\prime\prime},\,b=b^{\prime}+ib^{\prime\prime}$ and $A,B\in{\bf Mat}_{2}({\mathbb{R}})$ are defined as in (2.16), so that $(b_{\mathbb{R}}^{\theta})^{-1}\iota_{\theta}(u)b_{\mathbb{R}}^{\theta}=k_{\infty}(A+iB)$ (Lemma 2.3), then a computation reveals $C\bar{u}C^{-1}=A-iB$. Thus, the automorphism $u\mapsto C\bar{u}C^{-1}$ of ${\bf SU}(2)$ brings the subgroup ${\mathsf{B}}^{\#}({\mathbb{R}})\cap({\bf K}^{\#}_{\infty})^{0}=\{\left[\begin{smallmatrix}a&0\\ 0&a^{-1}\end{smallmatrix}\right]\mid a\in{\mathbb{C}}^{1}\}$ to the subgroup ${\bf SO}(2)$, by which $\mathsf{v}_{\varrho}^{0}$ is fixed. Hence, $\mathsf{f}_{\varrho}$ is left ${\mathsf{B}}^{\#}({\mathbb{R}})\cap({\bf K}^{\#}_{\infty})^{0}$-invariant. Thus, by $\mathsf{G}^{\#}({\mathbb{R}})={\mathsf{B}}^{\#}({\mathbb{R}})({\bf K}_{\infty}^{\#})^{0}$, there exists a unique element $\mathsf{f}^{(s)}_{\varrho}$ of ${\mathcal{V}}^{\#}(s,1,\mu_{\infty})$ such that $\mathsf{f}_{\varrho}^{(s)}|{\bf K}_{\infty}^{\#}=\mathsf{f}_{\varrho}$. Define $\phi_{\infty}\in{\mathcal{S}}({\mathbb{C}}^{2})$ by
@@
(4.24) $f_{\phi_{\infty}}^{(s,1,\mu_{\infty})}=|D|^{\frac{s+1}{2}}(-1)^{\frac{l_{1}-l_{2}}{2}}{\Gamma_{{\mathbb{C}}}\left(s+\tfrac{l_{1}-l_{2}}{2}+1\right)}\,\mathsf{f}_{\varrho}^{(s)}.$
@@
(4.25) $Z^{(\infty)}_{v}(\phi_{\infty},\varphi^{0}_{\pi};s,\mu_{\infty},\Lambda)$ $=w_{D}^{-1}\overline{R(\varphi^{0}_{\pi},E,\Lambda)}(v\mid v_{\varrho}^{0})_{\varrho}(-1)^{l_{1}}|D|^{\frac{1}{2}(1-\frac{l_{1}+l_{2}}{2})}{L\left(s+\tfrac{1}{2},\pi_{\infty}\right)}.$
The formula in [48, 7.23 Lemma] yields
@@
(4.27) $f_{\widehat{\phi_{\infty}}}^{(-s,1,\mu_{\infty})}=\tfrac{{|D|}}{4}(-1)^{\frac{l_{1}-l_{2}}{2}}f_{\phi_{\infty}}^{(-s,1,\mu_{\infty})}.$
@@ 4.6. The spinor $L$-function and its functional equation
Let $\pi\cong\otimes_{v}\pi_{v}$ be a cuspidal automorphic representation of $\mathsf{G}({\mathbb{A}})$ with the trivial central character: then, $\bar{\pi}_{p}\cong\pi_{p}^{\vee}\cong\pi_{p}$ for all $p<\infty$ by [45, Proposition 2.3]. The twist $\pi_{\mu}$ of $\pi$ by an idele class character $\mu:{\mathbb{A}}^{\times}/{\mathbb{Q}}^{\times}\longrightarrow{\mathbb{C}}^{1}$ is defined on the space $V_{\pi}$ of $\pi$ as $\pi_{\mu}(g):=\pi(g)\cdot(\mu\circ\nu)(g)$ for $g\in\mathsf{G}({\mathbb{A}})$, so that the central character of $\pi_{\mu}$ is $\mu^{2}$. Let $\pi\in\Pi_{\rm cusp}^{(T_{\theta},\Lambda)}(\lambda,N)$ and $\mu$ be as in §4.3, so that $\lambda=(l_{1},l_{2})\in{\mathbb{Z}}^{2}$, $l_{1}\equiv l_{2}\pmod{2}$, $l_{1}\geqslant l_{2}>2$ and the ramification loci of $\pi$ and $\mu$ are disjoint. We define the spinor $L$-function of $\pi$ twisted by $\mu$ as the Euler product
@@
with $L(s,(\pi_{\mu})_{p})$ the local $L$-factor listed in [36, Table A.8]. Define
@@
as expected. Moreover $\widehat{L}(s,\pi,\mu)$ is holomorphic except for possible simple poles at $s=3/2,-1/2$, which does not occur when $\mu$ is non-trivial. By (4.26), $M(s)\mathsf{f}_{\varrho}^{(s)}$ has a simple zero at $s=1$ if $l_{1}>l_{2}$, which in turn implies the holomorphy at $s=1$ of the global intertwining operator applied to the section $f_{\phi}^{(s)}$ as well as the Eisenstein series $E(\phi,s,\Lambda,\mu)$ for $\phi_{\infty}$ as above. Hence, by (4.20) with an appropriate $\phi$, $\widehat{L}(s,\pi,\mu)$ is holomorphic at $s=3/2$ when $l_{1}>l_{2}$. If $\mu$ is real valued, then $G(\tilde{\mu})/\sqrt{M}\in\{1,i\}$ and $\varepsilon(1/2,\pi,\mu)=(-1)^{l_{2}}$, so that $L(1/2,\pi,\mu)=0$ unless $l_{2}$ is even.